OpenStudy (anonymous):

how do you find the value of the an expression. x=-2/3 and y=-1/4

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

it says x/y

OpenStudy (anonymous):

k so a start is \[\large \frac{-\frac{2}{3}}{-\frac{1}{4}}\]

OpenStudy (anonymous):

how many minus signs?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

k so will the answer be positive or negative?

OpenStudy (anonymous):

a positive. I think we figured it out but I am not sure if it is right

OpenStudy (anonymous):

right, positive so we can ignore the annoying minus signs and start with \[\frac{2}{3}\times \frac{4}{1}\]

OpenStudy (anonymous):

now multiply straight across and you are done

OpenStudy (anonymous):

that's what we done, thank you so much. One more question the next one has only one minus sign how do you figure that out. Our math teacher doesn't help very much in class and my mom is right here helping but this is just refreshing her.

OpenStudy (anonymous):

k no problem if there is only one minus sign then the "final answer" will be negative

OpenStudy (anonymous):

ok thank you so much for your help!!!

OpenStudy (anonymous):

yw